# How Does a Deep Neural Network Look at Lexical Stress?

### v1

@@ IV.1 Layer-wise Relevance Propagation
In our study, we apply this technique (as implemented in the Captum library [14]) to our lexical stress classifiers, backpropagating prediction to time-frequency bins in the input. This method identifies the regions of the spectrogram that have the most significant influence on the model’s classification of lexical stress. We specifically examine our best-performing VGG-based classifier, VGG16, and ResNet18 (for model performance comparisons see Section V and Table 1).
Denote by $\mathbf{x}^{l}$ the input to the $l$-th layer. Denote by $\mathbf{W}^{l}$ the weights associated with layer $l$. Denote by $\mathbf{Z}^{l}$ is the elementwise preactivation of layer $l$ and it is defined as $Z_{ij}={W}^{l}_{ij}\odot{x}^{l}_{i}$. The relevance score of layer $l$ is computed recursively from a higher layer $l+1$. Relevance score $\mathbf{LRP_{Z}}$ is defined as [13]
@@
In this rule, a small constant $\epsilon$ is added to the denominator along with the sign of $Z^{l}_{j}$. This modification prevents division by zero and mitigates the impact of small or noisy preactivations $Z^{l}_{ij}$ on the relevance distribution. By incorporating $\epsilon$, the $\mathbf{LRP}_{\epsilon}$ rule enhances the stability of relevance propagation, leading to more reliable and interpretable results. This adjustment is particularly beneficial when dealing with minimal or problematic contributions, as outlined by Bach et al. [2].
The $\mathbf{LRP_{\alpha\beta}}$ rule, introduced by Bach et al. [2], is an extension of the basic LRP method that differentiates the contributions of facilitatory and inhibitory activation flow. It is defined as:
@@
Earlier implementations of LRP, such as those described in Bach et al. [2] and Lapuschkin et al. [16], typically employed a single LRP rule applied uniformly throughout the network. This approach failed to account for the varying needs of different layers or components within the network, resulting in less accurate and insightful relevance attributions. To address these limitations, Kohlbrenner et al. [13] introduced a composite strategy, which employs a combination of multiple LRP rules tailored to different parts of the network. In our analysis, an identity function was used to backpropagate activation for the first two convolutional layers closest to the input, $\mathbf{LRP}_{\alpha 1}$ ($\mathbf{LRP}_{\alpha\beta}$ with $\alpha=1$) for the remaining convolutional layers, and $\mathbf{LRP}{\epsilon}$ for the fully connected layers closest to the output. This combined approach, denoted as $\mathbf{LRP_{CMP}}$, enhances model interpretability and provides more accurate relevance attributions.
[…]
@@ IV.3 Feature-Specific Relevance
In addition to the IOU metric, we introduce a measure to evaluate how the model focuses on specific acoustic features during decision-making. Focusing on the vowel within the stressed syllable, this analysis aims to isolate the contribution of each acoustic feature to the model’s decision-making process. We constructed feature-specific heatmaps that represent the relevance distribution as if the model were focusing exclusively on the given feature. This was done by extracting the fundamental frequency ($F_{0}$), the first three formants ($F_{1},F_{2}$ and $F_{3}$) along with their formant bandwidth and sound intensity using the default settings of the Praat analysis software [3]. For each formant, heatmap values at each time point were generated by normalizing the intensity within the formant’s bandwidth and then scaling these normalized values by the sound intensity.
@@ V Model Performance
Classifier results on the test sets are summarized in Table 1. The VGG16 architecture achieved classification accuracies of 87% and 92% in distinguishing between initial and final stress in minimal pair and no minimal pair words test sets, respectively. The lower accuracy in the minimal pair test set may reflect errors in part-of-speech tagging. Note as well that human perception of lexical stress in disyllabic words based solely on acoustics (without access to syntactic or other contextual cues) is not at ceiling; [32] reported first-language English listeners had a mean accuracy of 84% in this task.
Other architectures also showed strong results. VGG19 matched VGG16’s performance, despite utilizing a more complex architecture, suggesting that VGG16 is sufficient. VGG11 showed slightly lower performance on both test sets. While the other architectures also showed slightly lower performance, they still achieved a high level of performance (>> 85% on both test sets).
Table 1: Classification accuracy for each model architecture on each test set.
@@
Minimal pair LeNet 85
VGG11 86
VGG16 87
VGG19 87
ResNet18 86
No minimal pair LeNet 90
VGG11 91
VGG16 92
VGG19 92
ResNet18 90
We also examined the ability of the classifier to generalize to other languages. While there is considerable cross-linguistic variation, there is also some degree of similarity in the acoustic realization of lexical stress across languages [4], leading us to expect some degree of generalization. We therefore examined the performance of the trained VGG16 and ResNet18 models before and after fine-tuning on both Dutch [24] and Hebrew [26] datasets. VGG16 achieved 93% accuracy on the Dutch data and 88% on the Hebrew data after fine-tuning. Similarly, ResNet18 achieved 93% on Dutch and an accuracy of 85% on Hebrew. This suggests our modeling framework provides a robust approach to detecting lexical stress. The results are summarized in Table 2.
Table 2: Classification accuracy for other languages.
Test Set Classifier Accuracy (%)
Dutch VGG16 81
ResNet18 79
VGG16 after fine-tuning 93
ResNet18 after fine-tuning 93
Hebrew VGG16 68
ResNet18 73
VGG16 after fine-tuning 88
ResNet18 after fine-tuning 85
@@ VI.1 Layer-wise Relevance Propagation (LRP) Analysis
Inspection of the LRP heatmaps for both initial and final stress words suggested they highlighted distinct areas of relevance that align with the energy distribution patterns observed in the spectrograms. Figures 2 and 3 illustrate this for the best-performing VGG16 classifier. For IS words, the LRP heatmaps emphasized the initial syllable as the most relevant area for classification by VGG16. Conversely, for FS words, the LRP heatmaps showed a concentration of relevance towards the end of the phoneme sequence. Figure 3 provides further illustrations of these differences, focusing on the $\mathbf{LRP_{CMP}}$ method.
@@ VI.2 Bounding Box and IOU Analysis
Table 4: $r$ and $p$ values for heatmaps based on different features in the stressed vowel region (averaged across samples), using VGG16 and applying $\mathbf{LRP_{CMP}}$ method (sorted by $r$; higher $r$ and lower $p$ are better). See text for details.
Method Acoustic Feature Mean $r$ Mean $p$
@@
VI.3 Feature-Specific Relevance Analysis
As shown in Table 4, a heatmap generated solely using $F_{1}$ of the stressed vowel exhibits the strongest correlation with the observed heatmap. Unsurprisingly, this formant makes a significant contribution, as it correlates with tongue height, jaw height, and the overall opening of the oral cavity, which in turn are strongly associated with vowel sound intensity. However, the combination of $F_{1}$ and $F_{2}$ is a close second to $F_{1}$ alone, suggesting that the two acoustic features traditionally related to vowel quality in English work together to determine classifier predictions. Similarly, heatmaps combining these two formants with $F_{3}$ and, to a lesser extent, $F_{0}$, are used to predict LRP responses. However, the outsize influence of $F_{1}$ is made clear by the low performance of the other formants and pitch in isolation.
@@ VI.3 Feature-Specific Relevance Analysis
What other acoustic features might explain these pixels? One possibility is that the classifiers are using this information to track pitch, a key contributor to lexical stress perception [4]. Our classifier relies on a standard spectrographic representation as input. The relatively poor resolution of this representation at low frequencies makes it a poor estimator of pitch. Tracking the spacing of harmonics (which have high amplitude in this region of the signal) would give the classifier access to this critical information. In fact, inspection of Figures 2 and 3 suggests that the classifier is attending to regularly spaced concentrations of energy in these lower frequencies – which could correspond to harmonics. This could allow the classifier to estimate pitch with a higher degree of resolution than is available in the region of the spectrogram corresponding to $F_{0}$, using information that can be readily uncovered from the spectrographic input.
@@ VII Discussion
To explore what allows CNNs to outperform more traditional approaches, we built several novel classifiers for lexical stress detection, using LRP-based analyses to explore the spatiotemporal properties that drive model performance. A fully automatic pipeline was used to extract disyllabic words from read and spontaneous speech. CNNs trained on words without minimal pairs exhibited high performance on held-out no minimal pair words and words with minimal pairs. Follow-up analyses showed that with fine-tuning, the classifiers exhibited high accuracy on two other languages (Dutch and Hebrew). LRP analyses on the spectrogram input space for the CNNs revealed that the models were sensitive to acoustic properties in both stressed and unstressed syllables, but showed greater reliance on stressed syllables (particularly the stressed vowel). Analysis of our highest-performing model showed that $F_{1}$ of the stressed vowel exerted a stronger influence on predictions than other formants, but $F_{2}$ and, to a lesser extent, $F_{3}$ also contributed. Furthermore, these features failed to capture a substantial amount of variance in the LRP heatmap, suggesting additional features of the stresseed vowel contribute to its predictions (e.g., spacing of harmonics, particularly below $F_{1}$, may be used to track pitch).
@@
The classifiers’ greater reliance on the properties of stressed versus unstressed syllables and more specifically stressed vowels versus other syllable regions is similar to the focus of the phonetics literature on stressed syllables and stressed vowels [4, 7]. Critically, differing from some early phonetics studies, the classifier does not focus exclusively on these regions, but also attends to information in both syllables, including non-vowel portions [^2] It is possible that distributed nature of stress cues reflects errors in our forced aligner. However, given the substantial temporal range of LRP pixels (see, e.g., Figure 4(c)), it’s likely that the results are robust to small changes to segment boundaries.. The use of information in the unstressed syllable resonates with the now-widespread use of relativized or normalized cues to lexical stress [4]. The use of information outside of the vowel is also consistent with phonetics research that examines whole-syllable duration and intensity [4]. More broadly, the use of consonantal information is consistent with the view that stress is not a property of vowels but of whole syllables, organized within more complex prosodic structures such as feet; these structres condition not only the properties of vowels but also surrounding consonants in stressed and unstressed syllables (see, e.g., [11]).
@@
Our feature-specific relevance analysis allowed us to connect the LRP results to formants of the stressed vowel. However, these analyses left a great deal of the LRP signal unexplained – not only in the stressed vowel, but also in other regions of the stressed syllable and the entirety of the unstressed syllable. Future work examining what acoustic features these pixels correspond to may provide new insights into the phonetics of English stress.

### v2

@@ IV.1 Layer-wise Relevance Propagation
In our study, we apply this technique [kokhlikyan2020captum] to our lexical stress classifiers, backpropagating prediction to time-frequency bins in the input [^3] We used the implemented in the Captum library https://captum.ai/api/lrp.html.. This method identifies the regions of the spectrogram that have the most significant influence on the model’s classification of lexical stress. We specifically examine our best-performing VGG-based classifier, VGG16, and ResNet18 (for model performance comparisons see Section V and Table 2).
Figure 3: Further illustrative examples of spectrograms and corresponding attributions for the words “subject,” “address” and “refuse” with different primary stress location, using the VGG16 model. Vertical lines and annotation follow Figure 2. Panel (a)-(d) show spectrograms and corresponding $\mathbf{LRP_{CMP}}$ heatmaps for the word ”subject” with initial and final stress, respectively. Similarly, panels (e)-(l) display spectrograms and heatmaps of the words “address” and “refuse,” with different stress types. Note that the examples were drawn from datasets in which word segments were zero-padded to reach the fixed 0.5 second window length, as described in Section II.2.
Denote by $\mathbf{x}^{l}$ the input to the $l$-th layer. Denote by $\mathbf{W}^{l}$ the weights associated with layer $l$. Denote by $\mathbf{Z}^{l}$ is the elementwise preactivation of layer $l$ and it is defined as $Z^{l}_{ij}={W}^{l}_{ij}\odot{x}^{l}_{i}$. The relevance score of layer $l$ is computed recursively from a higher layer $l+1$. In 9206975 the relevance score $\mathbf{LRP_{Z}}$ is defined as
@@
In this rule, a small constant $\epsilon$ is added to the denominator along with the sign of $Z^{l}_{j}$. This modification prevents division by zero and mitigates the impact of small or noisy preactivations $Z^{l}_{ij}$ on the relevance distribution. By incorporating $\epsilon$, the $\mathbf{LRP}_{\epsilon}$ rule enhances the stability of relevance propagation, leading to more reliable and interpretable results. This adjustment is particularly beneficial when dealing with minimal or problematic contributions, as outlined by bach2015pixel.
The $\mathbf{LRP_{\alpha\beta}}$ rule, introduced by bach2015pixel, is an extension of the basic LRP method that differentiates the contributions of facilitatory and inhibitory activation flow. It is defined as:
@@
Earlier implementations of LRP, such as those proposed by bach2015pixel and Lapuschkin2016CVPR, typically employed a single LRP rule applied uniformly throughout the network. This approach failed to account for the varying needs of different layers or components within the network, resulting in less accurate and insightful relevance attributions. To address these limitations, 9206975 introduced a composite strategy, which employs a combination of multiple LRP rules tailored to different parts of the network. In our analysis, an identity function was used to backpropagate activation for the first two convolutional layers closest to the input, $\mathbf{LRP}_{\alpha 1}$ ($\mathbf{LRP}_{\alpha\beta}$ with $\alpha=1$) for the remaining convolutional layers, and $\mathbf{LRP}{\epsilon}$ for the fully connected layers closest to the output. This combined approach, denoted as $\mathbf{LRP_{CMP}}$, enhances model interpretability and provides more accurate relevance attributions.
@@ IV.3 Feature-Specific Relevance
In addition to the IOU metric, we introduce a measure to evaluate how the model focuses on specific acoustic features during decision-making. Focusing on the vowel within the stressed syllable, this analysis aims to isolate the contribution of each acoustic feature to the model’s decision-making process. We constructed feature-specific heatmaps that represent the relevance distribution as if the model were focusing exclusively on the given feature. This was done by extracting the fundamental frequency ($F_{0}$), the first three formants ($F_{1},F_{2}$ and $F_{3}$) along with their formant bandwidth and sound intensity using the default settings of the Praat analysis software [pratt]. For each formant, heatmap values at each time point were generated by normalizing the intensity within the formant’s bandwidth and then scaling these normalized values by the sound intensity.
@@ V Model Performance
Classifier results on the test sets are summarized in Table 2. The VGG16 architecture achieved classification accuracies of 88% and 92% in distinguishing between initial and final stress in disyllabic minimal pair and no minimal pair words test sets, respectively. The lower accuracy in the minimal pair test set may reflect errors in part-of-speech tagging. Note as well that human perception of lexical stress in disyllabic words based solely on acoustics (without access to syntactic or other contextual cues) is not at ceiling; yu2010cross reported first-language English listeners had a mean accuracy of 84% in this task.
Other architectures also showed strong results. VGG19 matched VGG16’s performance, despite utilizing a more complex architecture, suggesting that VGG16 is sufficient. VGG11 showed slightly lower performance on both test sets. While the other architectures also showed slightly lower performance, they still achieved a high level of performance (>> 86% on both test sets).
Table 2: Classification accuracy for each model architecture on each stress type and test set.
@@
IS FS Overall
minimal pair LeNet 86.3 82.5 86
VGG11 87.2 83.5 87
VGG16 88.2 83.6 88
VGG19 88.3 82.3 88
ResNet18 87.74 84 87
no minimal pair LeNet 91 85.6 90
VGG11 91.8 88.3 91
VGG16 92.6 90.6 92
VGG19 92.2 91.3 92
ResNet18 90.3 88.7 90
@@ VI.1 Layer-wise Relevance Propagation (LRP) Analysis
Inspection of the LRP heatmaps for both initial and final stress words suggested they highlighted distinct areas of relevance that align with the energy distribution patterns observed in the spectrograms. Figures 2 and 3 illustrate this for the best-performing VGG16 classifier. For initial stress words, the LRP heatmaps emphasized the initial syllable as the most relevant area for classification by VGG16. Conversely, for final stress words, the LRP heatmaps showed a concentration of relevance towards the end of the phoneme sequence. Figure 3 provides further illustrations of these differences, focusing on the $\mathbf{LRP_{CMP}}$ method.
[…]
@@ VI.2 Bounding Box and IOU Analysis
VI.3 Feature-Specific Relevance Analysis
As shown in Table 4, a heatmap generated solely using $F_{1}$ of the stressed vowel exhibits the strongest correlation with the observed heatmap. Unsurprisingly, this formant makes a significant contribution, as it correlates with tongue height, jaw height, and the overall opening of the oral cavity, which in turn are strongly associated with vowel sound intensity. However, the combination of $F_{1}$ and $F_{2}$ is a close second to $F_{1}$ alone, suggesting that the two acoustic features traditionally related to vowel quality in English work together to determine classifier predictions. Similarly, heatmaps combining these two formants with $F_{3}$ and, to a lesser extent, $F_{0}$, are used to predict LRP responses. However, the outsize influence of $F_{1}$ is made clear by the low performance of the other formants and pitch in isolation. When additional features that are less relevant to the model’s actual decision-making are combined (e.g., $F_{1}$ + $F_{2}$ + $F_{3}$), the overall similarity with the LRP-derived heatmap tends to decrease. This reduction arises because the inclusion of non-informative or weakly correlated regions introduces additional variance, thereby lowering the $r$ value. In this sense, “more” does not necessarily mean “better”: although these additional features may still play supporting roles in classification, their weaker correspondence to the model’s main decision cues can dilute the overall correlation with the relevance pattern.
Table 4: $r$ and $p$ values for heatmaps based on different features in the stressed vowel region (averaged across samples), using VGG16 and applying $\mathbf{LRP_{CMP}}$ method, sorted by $r$. See text for details.
Method Acoustic Feature Mean $r\uparrow$ Mean $p\downarrow$
@@ VI.3 Feature-Specific Relevance Analysis
What other acoustic features might explain these pixels? Resolving questions like this is a key area for development of this work. The phonetics literature suggests a number of possibilities (which are not mutually exclusive). Intensity has been proposed as a cue to stress [cutler, gay1978physiological, gordon2017acoustic, vanHeuvenStress]. These pixels could reflect sensitivity to the concentration of energy in low frequencies during vowels. A number of prior studies have identified spectral tilt / spectral balance as an important cue to lexical stress [gordon2017acoustic, vanHeuvenStress]. Although spectral tilt measures often rely on information across a wider range of frequencies, spectral information in this range would provide a great deal of information about the steepness of spectral tilt. It is also possible that these pixels reflect sensitivity to harmonic structures. Inspection of Figures 2 and 3 suggests that the classifier is attending to regularly spaced concentrations of energy in these lower frequencies – which could correspond to harmonics. The classifiers are using this information to track pitch, a contributor to lexical stress perception [cutler]. Our classifier relies on a standard spectrographic representation as input. The relatively poor resolution of this representation at low frequencies makes it a poor estimator of pitch. Tracking the spacing of harmonics (which have high amplitude in this region of the signal) would give the classifier access to this critical information. This could allow the classifier to estimate pitch with a higher degree of resolution than is available in the region of the spectrogram corresponding to $F_{0}$, using information that can be readily uncovered from the spectrographic input. It is also possible that the classifier is using the first and second harmonic amplitude difference, H1-H2, to measure voice quality [garellek2019phonetics] differences related to stress [gordon2017acoustic].
@@
Figure 5: Classification accuracy for each minimal pair word obtained with the VGG16 model. Each bar represents the proportion of correctly classified tokens for a given minimal pair word. Bars corresponding to words with vowel reduction are shown in dark blue, and those without vowel reduction are shown in light blue with hatching. The number of tokens for each word is indicated in parentheses. Words are ordered by accuracy.
@@ VII Discussion
To explore what allows CNNs to outperform more traditional approaches, we built several novel classifiers for lexical stress detection in disyllabic words, using LRP-based analyses to explore the spatiotemporal properties that drive model performance. A fully automatic pipeline was used to extract disyllabic words from read and spontaneous speech. CNNs trained on words without minimal pairs exhibited high performance on held-out no minimal pair words and words with minimal pairs. LRP analyses on the spectrogram input space for the CNNs revealed that the models were sensitive to acoustic properties in both stressed and unstressed syllables, but showed greater reliance on stressed syllables (particularly the stressed vowel). Analysis of our highest-performing model showed that $F_{1}$ of the stressed vowel exerted a stronger influence on predictions than other formants, but $F_{2}$ and, to a lesser extent, $F_{3}$ also contributed. Furthermore, these features failed to capture a substantial amount of variance in the LRP heatmap, suggesting additional features of the stressed vowel contribute to its predictions (e.g., spectral balance and pitch based on spacing of harmonics).
@@
The classifiers’ greater reliance on the properties of stressed versus unstressed syllables and more specifically stressed vowels versus other syllable regions is similar to the focus of the phonetics literature on stressed syllables and stressed vowels [cutler, gay1978physiological]. Critically, differing from some early phonetics studies, the classifier does not focus exclusively on these regions, but also attends to information in both syllables, including non-vowel portions [^4] It is possible that distributed nature of stress cues reflects errors in our forced aligner. However, given the substantial temporal range of LRP pixels (see, e.g., Figure 4(c)), it’s likely that the results are robust to small changes to segment boundaries.. The use of information in the unstressed syllable resonates with the relatively widespread use of relativized or normalized cues to lexical stress [cutler, vanHeuvenStress]. The use of information outside of the vowel is also consistent with phonetics research that examines duration and intensity of the whole syllable, the rime, or consonants [cutler, gordon2017acoustic]. More broadly, the use of consonantal information is consistent with the view that stress is not a property of vowels but of whole syllables, organized within more complex prosodic structures such as feet; these structures condition not only the properties of vowels but also surrounding consonants in stressed and unstressed syllables (see, e.g., [jensen2000against]). Future work should continue to develop this general line of analysis by examining variation in stress driven by larger sentential and discourse contexts (see vanHeuvenStress for a review and discussion).
While stressed versus unstressed syllables are typically distinguished by full versus reduced vowel quality, some disyllabic English stress-pairs are pronounced with full vowels in both syllables. Figure 5 summarizes classification accuracy for each of the minimal pair words in the present study. Words whose unstressed syllable undergoes vowel reduction (e.g., PREsent vs. preSENT) are shown in dark blue, whereas words without vowel reduction (e.g., IMport vs. imPORT) are shown in light blue with hatching. Tagging of minimal pair words as “with” or “without” vowel reduction was based on a predefined list of English stress pairs [^5] The word list used to categorize minimal pairs as with and without vowel reduction is available at https://osf.io/5d4ks/overview. (n.b. the actual degree of reduction in particular productions may not precisely align with these assumptions). The figure reveals a clear performance gap: the models achieve substantially higher accuracy for words exhibiting vowel reduction than for those that do not. This pattern is consistent with the nature of the training data, which consisted primarily of no minimal pairs disyllabic words in which vowel reduction serves as an important acoustic correlate of stress [vanHeuvenStress]. It also supports the interpretation that formant information, particularly cues related to vowel reduction, plays a central role in model performance, while also confirming that the models are sensitive to additional cues in cases where vowel quality differences are minimal.
@@
Our feature-specific relevance analysis allowed us to connect the LRP results to formants of the stressed vowel. However, these analyses left a great deal of the LRP signal unexplained – not only in the stressed vowel, but also in other regions of the stressed syllable and the entirety of the unstressed syllable. This is consistent with prior work suggesting that formant information alone is insufficient for lexical stress classification [bentum2024processing]. Future work examining what acoustic features these pixels correspond to may provide new insights into the phonetics of English stress.